Train A and B are traveling in the same direction on parallel tracks. A is running 40 mph and Bis running 48 mph. A passes the station at 9:10am and B passes the station at 9:25am. What time will B catch up with A?
Answer by jorel1380(3719)

48n=40(n+1/4)
48n=40n+10
8n=10
n=10/8=5/4 hours=10:40
5/4(48)=60
5/4(40)+1/4(40)=50+10=60..